OpenStudy (anonymous):

Fred has 10 coins in a book bag. He has 4 quarters, 1 penny, and 5 dimes. If Fred reaches into his book bag and draws a coin without looking, what is the probability that he will draw a quarter? A. 10% B. 40% C. 25% D. 20%

OpenStudy (anonymous):

B. 40%

OpenStudy (anonymous):

how

OpenStudy (nurali):

Each coin has a 1/10 probability of being drawn. There are 4 quarters, therefore the probability of drawing a quarter is 4×1/10

OpenStudy (smored):

find the proportion of quarters from total. \[\frac{ 4 }{ 4+1+5 }=\frac{ 4 }{10 }=0.4=40percent\]

OpenStudy (anonymous):

okay... thank u.... smored can u write it a little faster

OpenStudy (smored):

sorry, takes me a bit of time for writing the equations.